OpenStudy (anonymous):

Give a unique example of each: An absolute value equation with two solutions An absolute value equation with one solution An absolute value equation with no solutions

OpenStudy (campbell_st):

| x + 3| = 11 ( 2solutions) you need to solve x + 3 = 11 and x + 3 = - 11 | x + 3 | = 0 ( 1 solution) you need to solve x + 3 = 0 |x + 3| = -10 ( no solutions) this is because absolute value will always return a value greater than or equal to zero...